OpenStudy (anonymous):

I Will Medal And Fan Best Answer! I need explanation AND answer! Kala drove 638 miles in 11 hours. At the same rate, how long would it take her to drive 464 miles?

OpenStudy (anonymous):

I'll take a look one sec

OpenStudy (anonymous):

This looks like a cross multiplication problem right?

OpenStudy (anonymous):

Well uh, anyway \[\frac{ 638 }{ 11 } = \frac{ 464 }{ ? }\] 464*11= 5104/638=8 so 8 hours?

OpenStudy (anonymous):

ty